how many grams p4 is formed for each mole of ca3(po4)2?


how many grammes of p4 is produced for each gram of ca3(po4)2 used?


how many moles sio2 is required for each mole of ca3(po4)2 used?


how many moles c is required for each mole of ca3(po4)2 used?

1
Expert's answer
2022-02-22T08:31:02-0500

2Ca3(PO4)2 "+" 6SiO2 "+" 10C "\to" P4 "+" 6CaSiO3 "+" 10CO


a)

Molar mass of P4 "=123.9"

Moles of P4 formed "=\frac{1}{2}"

Mass "=61.95g"



b)


Molar mass of calcium phosphate "=310.2"


Grams of P4 "=\frac{1}{2}×\frac{1}{310.2}×123.9=0.1997g"


c)

Moles of SiO2 "=\frac{1}{2}×6=3" moles


d)


Moles of C "=\frac{1}{2}×10=5" moles